OpenStudy (anonymous):

how many elements must a set have if the number of proper subsets of the set 1/2 of the total number of subsets of the set?

terenzreignz (terenzreignz):

If a set S has n-number of elements, then the TOTAL NUMBER OF subsets it will have is \[\huge 2^n\]However, one of these is the set S itself. Therefore, the total number of PROPER SUBSETS a set with n elements would have is \[\huge 2^n - 1\]

terenzreignz (terenzreignz):

And is this supposedly equal to \[\huge \frac12 \cdot 2^n\]

terenzreignz (terenzreignz):

So... \[\huge 2^n - 1 = \frac12 \cdot 2^n\]and solve for n.

OpenStudy (anonymous):

doing this equation will solve for n?

terenzreignz (terenzreignz):

Yes in effect. Pro-tip... let \(\large u = 2^n\) first, and solve for u.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

?

terenzreignz (terenzreignz):

Yes.

OpenStudy (anonymous):

so i would do 2^1-1=1/2*2^1?

terenzreignz (terenzreignz):

No, you've already solved for n... that's the number of elements your set has...

OpenStudy (anonymous):

ohhhh

terenzreignz (terenzreignz):

But yeah, you'll find that if you let n = 1, the equation checks out...

terenzreignz (terenzreignz):

It is in fact, the only value for n, which would make the equation true.